OpenStudy (anonymous):

What is the slope of the line between (3, -4) and (-2, 1)? 1 2 -2 -1

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ 3}}\quad ,&{\color{blue}{ -4}})\quad &({\color{red}{ -2}}\quad ,&{\color{blue}{ 1}}) \end{array} \\\quad \\ slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ y_2}}-{\color{blue}{ y_1}}}{{\color{red}{ x_2}}-{\color{red}{ x_1}}}\)

OpenStudy (compassionate):

\(\large Since~you~seem~new~here:\\\bf\Huge{~~\color{red}{\boxed{W}}\color{#FF9200}{\boxed{E}}\color{#FFDB00}{\boxed{L}}\color{#B6ff00}{\boxed{C}}\color{#00ff49}{\boxed{O}}\color{#00DBff}{\boxed{M}}\color{#B600ff}{\boxed{E}}\\~~~~~~~~~~~~~~~~~~\boxed{T}\boxed{O}\\~~~~~~~~~~~\color{#0092ff}{\boxed{O}}\color{#0092ff}{\boxed{P}}\color{#0092ff}{\boxed{E}}\color{#0092ff}{\boxed{N}}\color{#7cc517}{\boxed{S}}\color{#7cc517}{\boxed{T}}\color{#7cc517}{\boxed{U}}\color{#7cc517}{\boxed{D}}\color{#7cc517}{\boxed{Y}}\color{#7cc517}{\boxed{!}}} \\ \large\bf You~can~read~the~\underline{\href{ /code-of-conduct }{Code~of~Conduct}}~here. • \\ \it \large Now~let's~try~to~find~an~answer~to~your~problem~together.\)\ (3, -4) and (-2, 1) You want to use the slope formula. \[m = \frac{ y2 - y1 }{ x2 - x1 }\] Our points are: x2 = -2, x1 = 3 y2 = 1, y1 = -4 \[m = \frac{ 1 -(-4) }{ -2 - 3 } = \frac{ 1 + 4 }{ - 5 } = \frac{ 5 }{ -5 } = -1\] Thank you for using OpenStudy -- come back again! :)